2. Deponit is a nitroglycerin transdermal delivery system for the prevention of angina pectoris The in vivo release rate of the 16 cm2 patch is 0.2 mg hl. The reported volume of distribution is 3 L kg BW-1 and the observed clearance rate is 1 L kg l BW min 1. What is the steady state nitroglycerine concentration (nanograms per milliliter) in a 70 kg human? What is the elimination rate constant in 1 min i? What is the half-life of nitroglycerine in minutes
Answer #1

The volume of distribution is 3 L/kg/BW

Clearance rate is 1 L/kg/BW/min

Elimination rate constant =?

Half-life of nitroglycerin =?

Ke or elimination rate constant = clearance rate/volume of distribution

=> 1/3 = 0.33 per hour

=> 0.33/60 = 0.0055 per min.

Half-life of nitroglycerine is (0.693 * volume of distribution)/ clearance rate

=> (0.693 * 3)/ 1 = 2.079 minutes

It takes 2.079 minutes for nitroglycerin to become 50% in the plasma.

Bioavailability or fraction absorbed is 2 L/kg/BW/min (3-1)

Average steady state concentration of nitroglycerin is

C = (fraction absorbed*Dose)/(clearance *dosing interval)

=> (2*70 *2 *10^-7kg/hr )/ (1*70 * 60 mins)

=> 280*10^-7/4200

=> 6.66 * 10^-9 kg/L

=> 6.66 ng/ml (after conversion)